OpenStudy (zmudz):

If \(f(a + b) = f(a) + f(b) - 2f(ab)\) for all nonnegative integers \(a\) and \(b\), and \(f(1) = 1\), compute \(f(1986)\).

OpenStudy (irishboy123):

\[f(n) = f(1) + f(n-1) - 2f(1) f(n-1)\] \[ \implies f(n) + f(n-1) = 1\] f(2) = 0, f(3) = 1, f(4) = 0

OpenStudy (irishboy123):

and 1986 is even

ganeshie8 (ganeshie8):

that looks really neat!

OpenStudy (irishboy123):

is there proper mathese for this? ie a formal way to write it up?

ganeshie8 (ganeshie8):

to me it looks good the way it is now lets ask @freckles

OpenStudy (irishboy123):

great, thx!

OpenStudy (freckles):

I don't see a problem with what was said

ganeshie8 (ganeshie8):

what follows is really not necessary : \(f(n)+f(n-1)=1\) is a recurrence relation so we can solve it formally clearly a particular solution is \(f(n)=\frac{1}{2}\) for homogeneous solution, the characteristic equation is \(r+1=0\\\implies r=-1\) so the homogeneous solution is \(c*(-1)^n\) therefore the complete solution is given by \(f(n)=\frac{1}{2}+c*(-1)^n\)

ganeshie8 (ganeshie8):

\(c\) can be solved from the given initial condition, but i prefer just seeing the 1,0,1,0... pattern to the donkey work of solving the recurrence relation :)

OpenStudy (loser66):

@IrishBoy123 I don't get how you link the problem to f(n) = f(1) +.....

OpenStudy (irishboy123):

a + b = 1 + n - 1

OpenStudy (irishboy123):

that what you mean?

OpenStudy (loser66):

Got it, it is a smart interpretation.

OpenStudy (irishboy123):

i've tried googling "homogeneous solution, the characteristic equation " and all i'm getting are references to DE's.

OpenStudy (irishboy123):

ie i am trying to work out @ganeshie8 's post

OpenStudy (irishboy123):

honestly, just a steer fpr reading later, if there is one

ganeshie8 (ganeshie8):

he was letting \(a=1\) and \(b=n-1\) @Loser66 for sure i couldn't have figured out that substitution on my own

OpenStudy (loser66):

but the way he said n = n-1+1 is a perfect way to determine the function f(n). wwwwwwwwoah!!

OpenStudy (loser66):

I love this site, I learn the new thing everyday.

OpenStudy (loser66):

@ganeshie8 I don't think he set a =1, b = n-1

OpenStudy (loser66):

just f(n) in general and link it to f(1) then steer to a completely different from the original one. Solve the problem in general case, not just 1986 = 1+ 1985

OpenStudy (freckles):

http://openstudy.com/users/jagr2713#/updates/559d6afde4b0f93dd7c2b398 this problem reminds me of this one

ganeshie8 (ganeshie8):

at least thats how i interpreted his solution that induction idea seems interesting too

OpenStudy (freckles):

yeah I interpret that to a=1 and b=n-1

OpenStudy (loser66):

OMG, so I am the only one person who has the weirdest interpretation? hehehe... It's ok, I am crazy originally.

OpenStudy (freckles):

\[(f(a + b) = f(a) + f(b) - 2f(ab)) \\ a=1 \text{ and } b=n-1 \text{ gives } \\ f(1+(n-1))=f(1)+f(n-1)-2f(n-1)\]

ganeshie8 (ganeshie8):

@IrishBoy123 i think that pattern \(\large x_{complete} = x_{particular} + x_{homogeneous}\) is seen everywhere in math including linear algebra, number theory, differential equations etc...

ganeshie8 (ganeshie8):

as you know in linear algebra, to solve the system \(Ax=b\), the process is : 1) find a particular solution 2) find null solution then put the complete solution : \(\large x_{complete} = x_{particular} + x_{null}\)

ganeshie8 (ganeshie8):

in number theory, to solve a diophantine equation like \(2a+3b=7\), the process is : 1) find a particular solution 2) find null solution then put the complete solution : \(\large x_{complete} = x_{particular} + x_{null}\) (diophantine equation is an equation that needs to be solved over "integers")

ganeshie8 (ganeshie8):

you mentioned about differential eqns and we have seen recurrence relations in this thread already.. looks that pattern is common in any area of math that deals with solving equations..

OpenStudy (irishboy123):

ok got that! thx